Circular No. 8216

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 T1 (TRITTON) The predicted total magnitudes in the following ephemeris (H_{10} =

10.0), from the elements on IAUC 8215, are necessarily highly uncertain, as the comet is clearly in outburst when compared to its 1978 apparition.

2003TT α_{2000} δ_{2000} Δ r ϵ β Mag.
Sept. 28 8^{h}42.^{m}41 +21^{o}18.^{′}0 1.691 1.422 57.^{o}1 36.^{o}3 12.7
Oct. 3 8 58.26 +19 56.7 1.673 1.424 58.0 36.6 12.7
8 9 13.59 +18 31.3 1.657 1.429 59.1 36.9 12.6
13 9 28.37 +17 02.5 1.641 1.436 60.2 37.1 12.6
18 9 42.61 +15 31.1 1.626 1.445 61.4 37.3 12.7
23 9 56.28 +13 58.0 1.612 1.456 62.8 37.4 12.7
28 10 09.40 +12 23.8 1.598 1.469 64.2 37.5 12.7

S/2001 U 3 MPEC 2003-T29 has astrometry and orbital elements (a = 0.029 AU,

e = 0.14, i = 147^{o}, P = 267 days, H = 12.8) by B. G. Marsden for a new Uranian satellite observed on 2001 Aug. 13, 25, and Sept. 21 by M. Holman, B. Gladman, and collaborators (see IAUC 7980). Gladman identified further observations in 8.2-m UT3 Very Large Telescope images taken by P. Rousselot and O. Mousis on 2002 Sept. 3 and 5.

(66063) 1998 RO_{1}

P. Pravec, P. Kušnirák, and L. Šarounová, Ondřejov Observatory; P. Brown and G. Esquerdo, University of Western Ontario (UWO); D. Pray, Greene, RI, U.S.A.; L. A. M. Benner, Jet Propulsion Laboratory (JPL); M. C. Nolan, National Astronomy and Ionosphere Center; J. D. Giorgini and S. J. Ostro, JPL; and J.-L. Margot, California Institute of Technology, report that photometric observations — obtained during 2003 Sept. 16–27 at Ondřejov, at UWO, and by Pray — show the near-earth object (66063) to be a binary system with orbital period 14.53 hr. The primary rotates with a period of 2.492 hr, its low lightcurve amplitude indicative of a nearly spheroidal shape. The lightcurves suggest that the secondary has an aspect ratio of ∼ 1.5 and a rotation synchronous with its orbital period. Mutual eclipse/occultation events indicate a secondary-to-primary mean diameter ratio of ≥ 0.5. Arecibo (2380-MHz, 13-cm) delay-Doppler radar observations, obtained during Sept. 29–Oct. 4, show a spheroidal primary and an elongated secondary with separations at least as large as 600 m.

2003 October 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green